OpenStudy (anonymous):

WILL FAN MEDAL AND GIVE 2 OWL BUCKS :) A rectangular office is 14 feet by 10 ft. Both the length and the width are expanded by X feet so that the office is three times it original size. Which of the following equations can be used to find the value of x? A. 140=X^2 B.420=3x^2 C.140=140+24x+x^2 D.420=140+24x+x^2

OpenStudy (anonymous):

B

Directrix (directrix):

@clanger555 How did you get that?

OpenStudy (anonymous):

I guessed

OpenStudy (anonymous):

I don't understand how they got B.. /:

OpenStudy (anonymous):

I mean A not B

OpenStudy (anonymous):

It's Z Hope it helps :)

OpenStudy (anonymous):

Z?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

There is no Z /:

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

It's A because since it's being expanded by the length and width you will do x to the second power of 2 and the original area is 420 hope this helped

OpenStudy (anonymous):

No... I looked it up and it says it's E

OpenStudy (anonymous):

Hope it helped you! :D

OpenStudy (anonymous):

There is no E...

OpenStudy (anonymous):

OMG!!!!!!!!!!!! V?

OpenStudy (anonymous):

I'm done I'm like so done with this

Directrix (directrix):

the office is three times it original size. (14 + x) * (10 + x) = 3* (14)*(10) x^2 + 24x + 140 = 420 --> @finiteok Which option is this?